Status Report

NASA Space Station On-Orbit Status 2 Dec 2003

By SpaceRef Editor
December 2, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 2 Dec 2003

All ISS systems continue to function nominally, except as noted previously or below.

For CDR Michael Foale, it was time again for another general MedOps PFE (periodic fitness evaluation), checking up on blood pressure and electrocardiogram during programmed exercise on the CEVIS ergometer bike in the Lab. Readings were taken with the BP/ECG (blood pressure/electrocardiograph) by Kaleri, who assisted his crewmate as CMO (crew medical officer). (Last time done: 11/5).  [BP/ECG provides automated noninvasive systolic and diastolic blood pressure measurements while also monitoring and displaying accurate heart rates on a continual basis at rest and during exercise.]

Foale terminated the 24-hour charging process for EMU batteries #2029 & #2030 in the Airlock’s battery stowage assembly (BSA), then initiating their discharge after the required one-hour rest period. The SSC7 laptop needed for the procedure was first reconfigured for the discharge, then restored to its original settings. After the discharge, the batteries remain in the BSA. [The periodic charge/discharge cycle is a common maintenance requirement to restore the batteries’ capacity, prolonging their useful life.]

The FE performed the scheduled troubleshooting of the Russian VELO cycle ergometer (VB-3), checking, among else, the 5V power supply in the control unit in the cycle’s frame and testing the functionality of the generator assembly under large loading, followed by VELO reassembly. [For the testing, the BITS2-12 onboard telemetry measurement subsystem, in VD-SU control system mode, had to be connected to the VELO systems, which required prior deactivation of the SKV-1 air conditioner and Elektron O2 generator. After the testing, VD-SU and BITS telemetry connectors were remated and Elektron and SKV started up again.]

Alex Kaleri continued the inventory audit of comm network items begun by the Expedition 7 crew, for housekeeping and disposal of unneeded items. [The task was to sort out and list, by labeling and barcodes, all floppy and CD disks collected by Malenchenko and Lu in CTB (crew transfer bag) #1110. The list was then prepared for downlink via Regul Packet.]

The CDR completed his fifth weekly filling out of the Food Frequency Questionnaire (FFQ), which keeps a log of his nutritional intake over time on the MEC (medical equipment computer). [Last time done: 11/26.]

In the DC-1 “Pirs” docking module, Kaleri conducted his second regular (monthly) checkup of AZS circuit breakers on the BVP amp switch panel — they should all be On — and the 14 LEDs of the fuses in fuse panels BPP-30 and BPP-36 (last time done: 11/1).

Yesterday’s preparations for the Russian Plasma Crystal-3 (PK-3) payload by FE Alexander Kaleri were not completed, because the seal integrity of the vacuum chamber (ZB) could not be maintained. As a consequence, today’s planned PK-3 activities were aborted pending further leak checks and troubleshooting.  [The experiment involves zero-G crystallization processes in plasma, i.e., fine particles, charged and “excited” by RF/radio frequency power, inside the evacuated work chamber.]

Sasha did the daily SOZh life support systems maintenance (including toilet facility, food containers, water containers and solid waste containers) plus inspection of the BRPK air/condensate separator. He also prepared the daily IMS (inventory management system) update file.

Mike unstowed and set up the equipment for the FOOT (Foot/Ground Reaction Forces during Space Flight) experiment scheduled for tomorrow’s scheduled experiment run, preparing the equipment for EMG (electromyography) calibration with camcorder/video recording and the actual data take in the specially equipped outfit. Mike’s troubleshooting attempt yesterday to transfer his personal montage file to a flash card failed, but he “worked around” it by downloading the file to a floppy. [On 11/21, the CDR had performed a practice run of the calibration exercises necessary before actual FOOT activity.]

As a follow-up to yesterday’s S-band transmission test by MCC-H in low data rate (LDR), Kaleri today supported an audio communications test on both S-band S/G (space-to-ground) loops, by checking on the sound volume of a call from TsUP at various RS locations, via speakers of the SM’s STTS audio subsystem.

At 12:25pm EST, Foale and Kaleri supported a 20-min. televised PAO interview via Ku-band & S-band with Charles Osgood (for CBS’ “Osgood Files”), Bill Harwood and Peter King (for CBS News).
The crew worked out according to their regular daily physical exercise program of 2.5 hrs on the TVIS treadmill, RED exerciser, CEVIS cycle ergometer and VELO bike (with load trainer).

Both crewmembers conducted the weekly TVIS maintenance, with a time and date check done by the exercising crewmember just prior to powerdown or end of workout session.

TVIS Update: Passive TVIS exercise (gyro and stabilizers off) is currently performed, by crew choice, without TVIS handrail and with SPDs (subject positioning devices) in “outrigger” configuration (i.e., rotated away from the treadmill 90 degrees perpendicular), SPD hooks being attached to bungee cords connected between two handrails on the Service Module (SM) floor. SM handrails are deployed for use if needed for assistance. The crew is not required to reduce workout loads on the treadmill.  [In this config, exercise is allowed for two weeks, during which additional IWIS (internal wireless instrumentation system) and TVIS data need to be collected before the treadmill can be used beyond that time frame. The TVIS testing and data takes are scheduled for next Thursday and Friday (11/4-5). Crew inspection of the TVIS base, after skirt removal, has discovered that two of the corner stabilizer wire ropes are heavily frayed (“90%”). Downlinked video is being assessed.]

CMG Update: For continued monitoring of the U.S. CMGs (control moment gyroscopes), MCC-H will initiate tomorrow an on-board data “dump” to capture high-rate data for CMG-2, -3 and -4 (CMG-1 is failed).  [Downlink of the data dump is to be performed if vibration on any CMG reaches the value of 0.08 g (eight hundredth of a G), even for a single pulse (spike). If there is a LOS period (loss-of-signal, i.e., no-comm) within three hours of a spike exceeding 0.08g, the data recorded for that period on the ZOE (zone-of-exclusion) recorder are dumped. These procedures will remain in force until 12/9. The continued monitoring is attempting to characterize the 11/8 event of a previously unseen vibration signature on CMG-3.]

Starting this morning at 5:20am EST, TsUP/Moscow is running a 23h 57m test of the Russian ASN-2401 satellite navigation antenna system on the SM, which uses GLONASS/Uragan satellites (the Russian equivalent of GPS) to correct the on-board state vector (SV) information (i.e., ISS position and velocity referenced to an epoch, or time “hack”). [Up to now, SV corrections/updates are uplinked daily from the ground, along with time setting, time synchronization and various power bus checks, or they require SV transfers from the U.S. segment from time to time. The ASN equipment was originally installed in the SM but was found faulty and had to be returned to the ground. After repair it was shipped again to the station on Progress 11P, re-installed by Yuri Malenchenko on 7/8/03, and subjected by TsUP controllers to several months of testing, ending on 11/30.]

Planning is underway at TsUP for a major R&R (removal & replacement) task involving the BTA heat exchanger components of the Russian segment’s SKV-2 air conditioner.  [Changing the SKV-2 heat exchanger blocks was decided on because current inefficient performance of SKV-2 equates to lack of redundancy for the operating SKV-1. The R&R activity, performed similarly by Expedition 4’s Yuri Onufrienko and Carl Walz on SKV-1 on 4/3/2002, was originally planned for Increment 7 but was cancelled since the safety certificate for the necessary liquid Freon venting to space was not available (it is now). The R&R will probably be conducted on 12/15-12/26, split into four stages of one day each, with crew involvement in three of them. A crew review of the planned work will be scheduled beforehand (about 12/12).]

The cabin air repress from Progress 12P over the weekend raised the overall atmospheric pressure by approximately 10 mmHg. The MCA (major constituents analyzer) was successfully activated for the repress and returned to LEM (life extending mode) after its completion.

Yesterday’s crew search for the missing PBA EHTK (portable breathing assembly–extension hose/tee kit)did not locate the “lost” item. [A part of PEP (personal emergency provisions), the EHTK, which is assigned to the Node starboard emergency equipment locker, is required in the event of an emergency.]

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, wereIndustrialized SE Africa haze (good pass along the length of the subcontinental escarpment: looking left for plateau air mass quality and right for coastal plains air quality. Good views at nadir of “Africa’s Switzerland,” the small mountain kingdom of Lesotho), Delhi, India (nadir pass. The city is located where road and rail lines converge on the major Yamuna River), Cairo, Egypt (looking just right of track), Nile River Delta(nadir pass. Mapping swaths from margin to margin are requested for studies of land use change. Salt poisoning of soils is a major problem in the delta due to increased availability of water from Lake Nasser, allowing for three crops per year instead of the age-old one crop per year. The resulting increase in evaporation and lack of the annual flood to wash soils combine to cause salt build-up. The same acreage is lost within the delta due to salt poisoning as is gained by agricultural expansion into the desert beyond the delta. Both processes are well illustrated remotely), Alexandria, Egypt (looking a touch left on the coast), Tel Aviv, Israel (looking right on Israel’s coast), Bamako, Mali (good nadir pass. The city is hard to see: crew was to look on the east side of the main Niger River channel, where transport lines converge. Also, pointing the camera right and left for any fires in the surrounding Sahel),Puebla, Mexico(pointing right of track at the foot of the volcanic chain), andEl Paso, Texas(nadir pass).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 20A. Vozdukh CO2 scrubber is On (in Manual Mode 5/3). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. MCA (major constituents analyzer) is off (in Life Extending Mode). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 25.6; ppO2 (mmHg) — 157.7; ppCO2 (mmHg) — 5.2.
  • SM Transfer Compartment: Pressure (mmHg) — 761; temperature (deg C) — 19.3.
  • FGB Cabin: Pressure (mmHg) — 756; temperature (deg C) — 22.3.
  • Node: Pressure (mmHg) — 754.70; temperature (deg C) — 22.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 756.73; temperature (deg C) — 25.1; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 756.83; temperature (deg C) — 23.4; shell heater temp (deg C) — 23.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 22.2
  • PMA-2: Shell heater temp (deg C) — 20.0

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed (dual-angle) Position (non-suntracking). 
  • SM batteries: Battery #8, formerly known as #7, is still in slot #8 for troubleshooting; all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries: Battery #5 is off; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational (all lanes reintegrated 11/5).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; string #3 dropped out 10/22.

Propulsion System:

  • Total propellant load available:3704 kg (8166 lb) as of 11/13 [SM(755) + FGB(2597) + Progress M(352) + Progress M-1(0)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit (as of this morning, 8:11am EST [= epoch]):

  • Mean altitude — 372.8 km
  • Apogee — 376.9 km
  • Perigee — 368.7 km
  • Period — 92.0 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0006069
  • Orbits per 24-hr. day — 15.65
  • Mean altitude loss in last 24 hours — 200 m
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 28732
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.